# Computing in a Faulty Congested Clique

### v1

@@ Abstract
As an exemplifying application of our approach, we show that the $O(n^{1/3})$-round complexity of semi-ring matrix multiplication [Censor-Hillel, Kaski, Korhonen, Lenzen, Paz, Suomela, PODC 2015] remains the same up to polylog factors in the faulty model, even if the adversary can fail $99\%$ of the nodes.
@@ 1 Introduction
We consider deterministic algorithms, and study a faulty version of the $\mathsf{Clique~{}}$model, which we refer to as the $\mathsf{Faulty~{}Clique~{}}$model. In this model an adversary may fail nodes in the network throughout the computation, such that a failed node cannot continue to participate in the computation. Our goal is to construct a scheme that converts any $\mathsf{Clique~{}}$algorithm into a $\mathsf{Faulty~{}Clique~{}}$algorithm, with the smallest possible overhead to the round complexity of the algorithm. We capture the budget of failures that the adversary has as a parameter $c$, such that at least $n/c$ nodes must remain non-faulty. Thus, the adversary may fail $(\frac{c-1}{c})n$ nodes throughout the execution of an algorithm (see Section 2 for a formal definition of the model and the additional concepts used throughout the paper).
Without further restrictions on the adversary, it may fail a node at the very first round of the algorithm, preventing the others from ever completing the computation since they can never access the failed node’s input. Thus, the $\mathsf{Faulty~{}Clique~{}}$model must allow at least one initial quiet round, in which no faults can occur. We therefore consider the number of quiet rounds that are needed for a $\mathsf{Faulty~{}Clique~{}}$algorithm as an additional complexity measure which we aim to minimize. We emphasize that our method will require a very small constant number of quiet rounds. This does not render the model trivial for any task which can be solved in $O(1)$ rounds in the non-faulty $\mathsf{Clique~{}}$model, because if its exact complexity is larger than our constant then it cannot be entirely executed during the quiet rounds (also note that we do not modify the bandwidth by any constant, but rather we stick to a certain given bandwidth of $b\log n$ bits, as we cannot change the communication model based on the task that we wish to solve).
@@
Second, our main result improves upon this greatly for computing certain functions $f$: To this end, we identify two particular parameters which, informally, capture the locality of communication and the locality of computation that a certain computation requires, and show that our result obtains a super-fast computation of functions for which these parameters are not large (see Section 2 for formal definitions). For such functions, the complexity overhead of our approach can be as low as $\tilde{O}(1)$ for a constant $c$. An example for such a function is computing the semi-ring product of two matrices, whose complexity in the $\mathsf{Clique~{}}$model is $O(n^{1/3})$ due to [CKK+19], for which our approach implies a $\tilde{O}(n^{1/3})$-round $\mathsf{Faulty~{}Clique~{}}$algorithm.
A remark on quiet rounds. Since our approach will encode the input data at the very start of the algorithm during the quiet rounds, we mention that it seems tempting to simply define the model as one in which the inputs are given in an encoded manner to the nodes to begin with. This also seems aligned with our requirement from the outputs. However, it turns out that the manner in which the inputs are encoded affects the overall complexity (see end of Subsection 1.1), thus it becomes necessary to carefully describe (and measure the complexity thereof) the initial quiet rounds.
@@ 1.1 Our Contribution
Our treatment of $\mathsf{Clique~{}}$and $\mathsf{Faulty~{}Clique~{}}$algorithms goes through circuits. This is aligned with prior work of [DKO14], which show how to compute circuits with $\mathsf{Clique~{}}$algorithms. For completeness and for setting the ground with the terminology that we need, we define in Section 2 the concept of layered circuits. Roughly speaking, these are circuits whose gates can be nicely split into layers, with wires only between two consecutive layers. When we compute a circuit by a $\mathsf{Clique~{}}$or $\mathsf{Faulty~{}Clique~{}}$algorithm, we assign each node with the task of computing some of the gates in a layer, which we refer to as this node’s part.
Layered circuits can be parametrized by what we call their parallel partition parameter. Informally, this parameter relates to the number of gates that have wires into and from the gates of any node’s part. This captures the amount of “communication” between layers when we later interpret them as $\mathsf{Clique~{}}$algorithms. Intuitively, we say that a layered circuit has a parallel partition of $n^{\varphi}$ when its gates have wires into/from $O(n^{1+\varphi})$ gates in other parts. Thus, when we compute a layer of the circuit by a $\mathsf{Clique~{}}$algorithm, the multiplicative overhead beyond the $n$ elements that each node can send/receive in a round is $O(n^{\varphi})$. This means that the $\mathsf{Clique~{}}$algorithm will incur this number of rounds per layer that it computes. It is not hard to show that any $\mathsf{Clique~{}}$algorithm that runs in $T$ rounds has a circuit of depth $O(T)$ with a 1-parallel partition ($\varphi=0$). We do so in Lemma 3 in Section 3.
@@
In particular, Theorem 2 means that for any constant $c$, we incur only a logarithmic overhead in the round complexity over the non-faulty $\mathsf{Clique~{}}$model, in which this task can be solved in $O(n^{1/3})$ rounds due to [CKK+19]. We emphasize that matrix multiplication is just an example, and there are more cornerstone tasks that can enjoy our approach by constructing circuits that are efficient for the use of Theorem 1.
@@
Another remark on quiet rounds. Matrix multiplication is an excellent example for why we cannot simply replace quiet rounds by an assumption of coded inputs. If we do not carefully shuffle the input matrices before encoding them, then the nodes will be forced to download (decode) too much information compared to what they need, resulting in communication and computation localities that are extremely higher compared to those of the circuit that we construct for proving Theorem 2. More explicitly, if we want a matrix multiplication algorithm to re-distribute the entries of the second input matrix among the nodes such that each node holds a column of the matrix rather than a row, or if we want each node to hold any other pattern of $n$ entries as we do in Section 5, then we need the information to already be encoded in this manner, as otherwise a node has to decode too much information in order to obtain its relevant entries. This could result in a huge number of rounds for a task that takes $O(1)$ rounds in the non-faulty $\mathsf{Clique~{}}$model.
Leveraging a sublinear bound on the number of faults. Finally, we show that if the bound on the number of failures is sublinear, then our method allows circuits with relaxed properties to be implemented by fast $\mathsf{Faulty~{}Clique~{}}$algorithms. The high level intuition is that if we are promised no more than $(c-1)n^{\chi}/c$ faults, we can replace the usage of codewords of length $n$ with $n^{1-\chi}$ disjoint codewords of length $n^{\chi}$. Since these are only technical modifications, we defer them to Section 6. This allows us to further extend our results for fast (ring) matrix multiplication, for which we show an algorithm with a round complexity of $O(c^{2}\log{n}\max\{n^{\chi},n^{1-2/\omega}\})$, where $\omega$ is the exponent of matrix multiplication (and $1-2/\omega$ is the exponent of $\mathsf{Clique~{}}$matrix multiplication [CKK+19]). See Section 7.
@@ 1.2 Technical Overview
To prove Theorem 1, our approach is as follows. First, we partition the gates of each layer of the circuit into its respective parts in the parallel partition, and assign one node to each part. In the first $O(1)$ quiet rounds, we shuffle the inputs of the nodes such that each node holds the input wires to its gates, using Lenzen’s routing scheme [Len13]. Then, each node $w$ encodes its state to prepare for the possibility of a failure, and sends one piece of the codeword to every other node. The code that is used has to be able to tolerate $(\frac{c-1}{c})n$ erasures, as this is the number of nodes that may fail (a node that has already failed simply does not receive its piece from $w$). We use $[n,n/c,(\frac{c-1}{c})n+1]_{q}$ codes, where $q$ is a field whose elements are of $\Theta(cb\log{n})$ bits, and thus we need $c=O(1)$ additional quiet rounds for this step.
@@
We mention that, informally speaking, coding is essential for $\mathsf{Faulty~{}Clique~{}}$algorithms. We show in Lemma 7 that if the messages of the quiet rounds depend only on the input of each node and not on messages it receives, then the messages of $c$ quiet rounds must form codewords of an $[n,n/c,(\frac{c-1}{c})n+1]_{q}$ error-correcting code.
@@ 2.1 The Model
The $c$-$\mathsf{Faulty~{}Clique~{}}$model is similar to the $\mathsf{Clique~{}}$model, except that an adversary may fail up to a fraction of $(c-1)/c$ of the nodes throughout the execution of an algorithm (that is, $n/c$ nodes must be non-faulty). A node that is failed in a certain round, does not send or receive any message from that round on. The $\mathsf{Faulty~{}Clique~{}}$model allows an initial number of quiet rounds, in which no faults can occur.
@@
For a round $T_{\textrm{non-faulty}}$ of $\mathcal{E}_{\textrm{non-faulty}}$, we say that a possibly faulty execution $\mathcal{E}$ of $\mathcal{A}$ satisfies at round $T$ the $R$-decodability condition if the following holds. Suppose every node $u$ is associated with any other node $w_{u}$, then it is possible in $R$ rounds for each node $u$ that is non-faulty at the end of these $R$ rounds to obtain $\sigma_{w_{u},T_{\textrm{non-faulty}}}$. (Informally the only way the adversary can prevent $u$ from receiving knowledge of any other node’s state after the next $R$ rounds is by failing $u$ itself.)
@@ 2.2 Layered Circuits
For a node $v$ in a directed graph $G$, let $\mathrm{in}(v):=\{w\in V\mid(w,v)\in E\}$ and $\mathrm{out}(v):=\{w\in V\mid(v,w)\in E\}$. The values $|\mathrm{in}(v)|$ and $|\mathrm{out}(v)|$ are called the fan-in and fan-out of $v$, respectively.
@@
We say that ${G}$ has a parallel partition of size $n$ with block fan size $n^{1+\varphi}$ if there exists such a pair of partitions in which for every $i,j\in[n]$, $|E_{i}^{L}|,|E_{j}^{R}|={O(n^{1+\varphi})}$. We say that a layered circuit $\mathcal{C}=(V,E)$ of depth $n^{\delta}$ has an $n^{\varphi}$-parallel partition, if there is a refinement $\mathcal{P}$ of the partition of $V=V_{0}\cup V_{1}\cup...\cup V_{n^{\delta}}$ such that for every $i$, the restriction of ${\mathcal{P}}$ to each of the subgraphs $V_{i}\cup V_{i+1}$ is a parallel partition of size $n$ with block fan size $n^{1+\varphi}$ with $L=V_{i}$, $R=V_{i+1}$.
@@ 3 $\mathsf{Clique~{}}$Algorithms as Layered Circuits
Let $\mathcal{A}$ be a $\mathsf{Clique~{}}$algorithm that computes a function $f$ in $T$ rounds. There is a circuit $\mathcal{C}$ of depth $12T+$ 1 with alphabet size $|\Sigma|=b\log(n)$ that has a $1$-parallel partition (that is, $n^{\varphi}$ with $\varphi=0$), for which $f^{\mathcal{C}}=f$.
@@
Figure 1: Illustration for layers $\ell=0,\dots,3$. The gates of node 00 appear on the leftmost box, while for nodes $1$ and $n-1$ only several of their gates are depicted in the other two boxes. The solid thin arrows represent wires (some wires are omitted for clarity). A dashed wide arrow between two blocks of gates means that every gate in the block at the lower level has a wire into every gate in the block of the upper level.
Circuit parameters: By construction, the depth of the circuit is thus $12T+$ 1. In addition, every consecutive pair of gate layers has a parallel partition with block fan size of 1. To see this, consider $V_{\ell},V_{\ell+1}$. If $\ell=2\ell^{\prime}+1$ then for every $i\in[n]$, the only wires into gates of $P_{\ell+1,i}$ are from gates in $P_{\ell,i}$ and thus the block fan size is 0. Otherwise, if $\ell=2\ell^{\prime}$ then for every $i\in[n]$, the only wires into $P_{\ell+1,i}$ which are from gates in $P_{\ell,i^{\prime}}$ for $i^{\prime}\neq i$ are the wires $e^{(\ell-1)}_{j,i}:=v^{(\ell-1)}_{j,i}\rightarrow v^{(\ell)}_{i,j}$ for $j\in[n]$. Thus, each block has one incoming wire from every other block, so the block fan size is $n$ (i.e., $\varphi=0$).
@@
Induction Step: We assume the induction hypothesis holds up to layer $12\ell^{\prime}-$ 1 and prove it for layers $2\ell^{\prime}$ and $12\ell^{\prime}+$ 1.
@@
For layer $12\ell^{\prime}+$ 1, by construction, each gate $v^{(2\ell^{\prime}+1)}_{i,j}$ for $i,j\in[n]$ receives as input wires all outputs of gates $v^{(2\ell^{\prime})}_{i,k}$, for $k\in[(\ell^{\prime}+1)n]$. Since these are all inputs and messages that node $i$ receives, this gate correctly computes the message sent by node $i$ to node $j$ in round $\ell^{\prime}+1$ of $\mathcal{A}$. Finally, by construction, gate $v^{(2\ell^{\prime}+1)}_{i,n+j}$ for $j\in[n,(\ell^{\prime}+1)n]$ gets and outputs the wire from gate $v^{(2\ell^{\prime})}_{i,j}$ and thus corresponds to the input of node $i$ and all messages it received in rounds 1 to $\ell^{\prime}$, as needed. ∎

### v2

@@ Abstract
As an exemplifying application of our approach, we show that the $O(n^{1/3})$-round complexity of semi-ring matrix multiplication [Censor-Hillel, Kaski, Korhonen, Lenzen, Paz, Suomela, PODC 2015] remains the same up to polylog factors in the faulty model, even if the adversary can fail $99\%$ of the nodes (or any other constant fraction).
@@ 1 Introduction
We consider deterministic algorithms, and study a faulty version of the $\mathsf{Clique~}$model, which we refer to as the $\mathsf{Faulty~Clique~}$model. In this model an adversary may fail nodes in the network throughout the computation, such that a failed node cannot continue to participate in the computation. We capture the budget of failures that the adversary has as a parameter $c$, such that at least $n/c$ nodes must remain non-faulty. Thus, the adversary may fail $(\frac{c-1}{c})n$ nodes throughout the execution of an algorithm (see Section 2 for a formal definition of the model and the additional concepts used throughout the paper). We ask:
How efficiently can one compute in the $\mathsf{Faulty~Clique~}$model?
Our contribution is a scheme that converts any $\mathsf{Clique~}$algorithm into a $\mathsf{Faulty~Clique~}$algorithm. The goal of such a scheme is to incur the smallest possible overhead to the round complexity of the algorithm. We exemplify the strength of our scheme by showing that it allows multiplying two matrices over a semi-ring in $\tilde{O}(n^{1/3})$ rounds, thus retaining its non-faulty complexity from [CKK+19] up to polylog factors. This is more than quadratically faster compared to the complexity of $n^{2/3+o(1)}$ that can be obtained for this problem from the follow-up work of [CHFGS25].
Let’s dive into some background.
Without further restrictions on the adversary, it may fail a node at the very first round of the algorithm, preventing the others from ever completing the computation since they can never access the failed node’s input. Note that we do not allow loss of information since our requirement is that the output depends on the original inputs, as opposed to some other fault models that allow the output to depend only on the inputs of non-faulty nodes. Thus, the $\mathsf{Faulty~Clique~}$model must allow at least one initial quiet round, in which no faults can occur. We therefore consider the number of quiet rounds that are needed for a $\mathsf{Faulty~Clique~}$algorithm as an additional complexity measure which we aim to minimize. We emphasize that our method will require a very small constant number of quiet rounds. This does not render the model trivial for any task which can be solved in $O(1)$ rounds in the non-faulty $\mathsf{Clique~}$model, because if its exact complexity is larger than our constant then it cannot be entirely executed during the quiet rounds (also note that we do not modify the bandwidth by any constant, but rather we stick to a certain given bandwidth of $b\log n$ bits, as we cannot change the communication model based on the task that we wish to solve).111[CHFGS25] work with already-coded inputs and thus avoid the notion of quiet rounds. Here, we care about how the input is coded and hence we need these constant number of preprocessing rounds.
@@
Second, our main result improves upon this greatly for computing certain functions $f$: To this end, we identify two particular parameters which, informally, capture the locality of communication and the locality of computation that a certain computation requires, and show that our result obtains a super-fast computation of functions for which these parameters are not large (see Section 2 for formal definitions). For such functions, the complexity overhead of our approach can be as low as $\tilde{O}(1)$ for a constant $c$. This insight is what allows us to compute the semi-ring product of two matrices, whose complexity in the $\mathsf{Clique~}$model is $O(n^{1/3})$ due to [CKK+19], in $\tilde{O}(n^{1/3})$-rounds in the $\mathsf{Faulty~Clique~}$model.
@@ 1.1 Our Contribution
Our treatment of $\mathsf{Clique~}$and $\mathsf{Faulty~Clique~}$algorithms goes through circuits, as was also done in the follow-up work of [CHFGS25], and is aligned with prior work of [DKO14] which show how to compute circuits with $\mathsf{Clique~}$algorithms. For completeness and for setting the ground with the terminology that we need, we define in Section 2 the concept of layered circuits. Roughly speaking, these are circuits whose gates can be nicely split into layers, with wires only between two consecutive layers. When we compute a circuit by a $\mathsf{Clique~}$or $\mathsf{Faulty~Clique~}$algorithm, we assign each node with the task of computing some of the gates in a layer, which we refer to as this node’s part. Note that a node always refers to a computing component in the $\mathsf{Clique~}$or $\mathsf{Faulty~Clique~}$model, while we use gates and wires to refer to circuits.
Layered circuits can be parametrized by what we call their parallel partition parameter. Informally, this parameter relates to the number of gates that have wires into and from the gates of any node’s part. This captures the amount of “communication” between layers when we later interpret them as $\mathsf{Clique~}$algorithms. Intuitively, we say that a layered circuit has a parallel partition of $n^{\varphi}$ when its gates have wires into/from $O(n^{1+\varphi})$ gates in other parts. Thus, when we compute a layer of the circuit by a $\mathsf{Clique~}$algorithm, the multiplicative overhead beyond the $n$ elements that each node can send/receive in a round is $O(n^{\varphi})$. This means that the $\mathsf{Clique~}$algorithm will incur this number of rounds per layer that it computes. It is not hard to show that any $\mathsf{Clique~}$algorithm that runs in $T$ rounds has a circuit of depth $O(T)$ with a 1-parallel partition ($\varphi=0$). Notably, this does not depend on any non-adaptivity assumption on the $\mathsf{Clique~}$algorithm (that is, this applies also to algorithms in which the communication pattern may depend on the inputs). We show this in Lemma 4 in Section 3.
@@
In particular, Theorem 2 means that for any constant $c$, we incur only a logarithmic overhead in the round complexity over the non-faulty $\mathsf{Clique~}$model, in which this task can be solved in $O(n^{1/3})$ rounds due to [CKK+19]. Following up on our work, [CHFGS25] present a different scheme, which yields a complexity of $T^{2}\cdot n^{o(1)}$ rounds in the $\mathsf{Faulty~Clique~}$model for any $T$-round $\mathsf{Clique~}$algorithm. While for some $\mathsf{Clique~}$algorithms their approach improves upon ours, for others our approach is significantly faster: their scheme would give a complexity of $n^{2/3+o(1)}$ for semi-ring matrix multiplication, which is quadratically slower compared to ours. We emphasize that matrix multiplication is just an example, and there are more cornerstone tasks that can enjoy our approach by constructing circuits that are efficient for the use of Theorem 1.
@@
A remark on quiet rounds. Matrix multiplication is an excellent example for why we cannot simply replace quiet rounds by an assumption of coded inputs. If we do not carefully shuffle the input matrices before encoding them, then the nodes will be forced to download (decode) too much information compared to what they need, resulting in communication and computation localities that are extremely higher compared to those of the circuit that we construct for proving Theorem 2. More explicitly, if we want a matrix multiplication algorithm to re-distribute the entries of the second input matrix among the nodes such that each node holds a column of the matrix rather than a row, or if we want each node to hold any other pattern of $n$ entries as we do in Section 5, then we need the information to already be encoded in this manner, as otherwise a node has to decode too much information in order to obtain its relevant entries. This could result in a huge number of rounds for a task that takes $O(1)$ rounds in the non-faulty $\mathsf{Clique~}$model.
Regimes of lower number of faults. Finally, we show that if the bound on the number of failures is sublinear, then our method allows circuits with relaxed properties to be implemented by fast $\mathsf{Faulty~Clique~}$algorithms. The high level intuition is that if we are promised no more than $(c-1)n^{\chi}/c$ faults, we can replace the usage of codewords of length $n$ with $n^{1-\chi}$ disjoint codewords of length $n^{\chi}$. Since these are only technical modifications, we defer them to Section 6. This allows us to further extend our results for fast (ring) matrix multiplication, for which we show the following (see Section 7).
Suppose that the inputs on the nodes are matrices $A,B\in\Sigma^{n\times n}$, where each node begins with $n$ coefficients of $A$ and $B$ and $\Sigma$ is a ring, then the value $C=A\cdot B$ can be computed in the ($\chi,c$)-$\mathsf{Faulty~Clique~}$with $O(1)$ quiet rounds, a round complexity of $O(\max\{c^{2}n^{\chi}\log n,c^{2}n^{1-2/\omega}\log n\})$, and $O(1)$-decodability.
Here $\omega$ is the exponent of matrix multiplication and $O(n^{1-2/\omega})$ is the complexity of ring matrix multiplication in the non-faulty setting [CGH19].
@@ 1.2 Technical Overview
To prove Theorem 1, our approach is as follows. First, we partition the gates of each layer of the circuit into its respective parts in the parallel partition, and assign one node to each part. In the first $O(1)$ quiet rounds, we shuffle the inputs of the nodes such that each node holds the input wires to its gates, using Lenzen’s routing scheme [Len13]. Then, each node $w$ encodes its state to prepare for the possibility of a failure, and sends one piece of the codeword to every other node. The code that is used has to be able to tolerate $(\frac{c-1}{c})n$ erasures, as this is the number of nodes that may fail (a node that has already failed simply does not receive its piece from $w$). We use $n,n/c,(\frac{c-1}{c})n+1]_{q}$ codes, where $q$ is a field whose elements are of $\Theta(cb\log{n})$ bits, and thus we need $c=O(1)$ additional quiet rounds for this step.
@@
We mention that, informally speaking, coding is essential for $\mathsf{Faulty~Clique~}$algorithms. We show in Lemma 8 that if the messages of the quiet rounds depend only on the input of each node and not on messages it receives, then the messages of $c$ quiet rounds must form codewords of an $n,n/c,(\frac{c-1}{c})n+1]_{q}$ error-correcting code.
@@ 2.1 The Model
The $c$-$\mathsf{Faulty~Clique~}$model is similar to the $\mathsf{Clique~}$model, except that an adversary may fail up to a fraction of (c1)/c(c-1)/c of the nodes throughout the execution of an algorithm (that is, $n/c$ nodes must be non-faulty). A node that is failed in a certain round, does not send or receive any message from that round on. The $\mathsf{Faulty~Clique~}$model allows an initial number of quiet rounds, in which no faults can occur.
@@
For a round $T_{\textrm{non-faulty}}$ of $\mathcal{E}_{\textrm{non-faulty}}$, we say that a possibly faulty execution $\mathcal{E}$ of $\mathcal{A}$ satisfies at round $T$ the $R$-decodability condition if the following holds. Suppose every node $u$ is associated with some other node $w_{u}$, then it is possible in $R$ rounds for each node $u$ that is non-faulty at the end of these $R$ rounds to obtain $\sigma_{w_{u},T_{\textrm{non-faulty}}}$. (Informally, the only way the adversary can prevent $u$ from receiving knowledge of any other node’s state after the next $R$ rounds is by failing $u$ itself.)
@@ 2.2 Layered Circuits
For a vertex $v$ in a directed graph $G$, let $\mathrm{in}(v):=\{w\in V\mid(w,v)\in E\}$ and $\mathrm{out}(v):=\{w\in V\mid(v,w)\in E\}$. The values $|\mathrm{in}(v)|$ and $|\mathrm{out}(v)|$ are called the fan-in and fan-out of $v$, respectively.
@@
We say that G{G} has a parallel partition of size $n$ with block fan size $n^{1+\varphi}$ if there exists such a pair of partitions in which for every $i,j\in[n]$, $|E_{i}^{L}|,|E_{j}^{R}|={O(n^{1+\varphi})}$. We say that a layered circuit $\mathcal{C}=(V,E)$ of depth $n^{\delta}$ has an $n^{\varphi}$-parallel partition, if there is a refinement $\mathcal{P}$ of the partition of $V=V_{0}\cup V_{1}\cup...\cup V_{n^{\delta}}$ such that for every $i$, the restriction of $\mathcal{P}$} to each of the subgraphs $V_{i}\cup V_{i+1}$ is a parallel partition of size $n$ with block fan size $n^{1+\varphi}$ with $L=V_{i}$, $R=V_{i+1}$.
@@ 3 $\mathsf{Clique~}$Algorithms as Layered Circuits
Let $\mathcal{A}$ be a $\mathsf{Clique~}$algorithm that computes a function $f$ in $T$ rounds. There is a circuit $\mathcal{C}$ of depth $2T+1$ with alphabet size $|\Sigma|=b\log(n)$ that has a $1$-parallel partition (that is, $n^{\varphi}$ with $\varphi=0$), for which $f^{\mathcal{C}}=f$.
@@
Figure 1: Illustration for layers $\ell=0,\dots,3$. The gates of node 0 appear on the leftmost box, while for nodes $1$ and n1n-1 only several of their gates are depicted in the other two boxes. The solid thin arrows represent wires (some wires are omitted for clarity). A dashed wide arrow between two blocks of gates means that every gate in the block at the lower level has a wire into every gate in the block of the upper level.
Circuit parameters: By construction, the depth of the circuit is thus $2T+1$. In addition, every consecutive pair of gate layers has a parallel partition with block fan size of 1. To see this, consider $V_{\ell},V_{\ell+1}$. If $\ell=2\ell^{\prime}+1$ then for every $i\in[n]$, the only wires into gates of $P_{\ell+1,i}$ are from gates in $P_{\ell,i}$ and thus the block fan size is 0. Otherwise, if $\ell=2\ell^{\prime}$ then for every $i\in[n]$, the only wires into $P_{\ell+1,i}$ which are from gates in $P_{\ell,i^{\prime}}$ for $i^{\prime}\neq i$ are the wires $e^{(\ell-1)}_{j,i}:=v^{(\ell-1)}_{j,i}\rightarrow v^{(\ell)}_{i,j}$ for $j\in[n]$. Thus, each block has one incoming wire from every other block, so the block fan size is $n$ (i.e., $\varphi=0$).
@@
Induction Step: We assume the induction hypothesis holds up to layer $2\ell^{\prime}-1$ and prove it for layers $2\ell^{\prime}$ and $2\ell^{\prime}+1$.
@@
For layer $2\ell^{\prime}+1$, by construction, each gate $v^{(2\ell^{\prime}+1)}_{i,j}$ for $i,j\in[n]$ receives as input wires all outputs of gates $v^{(2\ell^{\prime})}_{i,k}$, for $k\in[(\ell^{\prime}+1)n]$. Since these are all inputs and messages that node $i$ receives, this gate correctly computes the message sent by node $i$ to node $j$ in round $\ell^{\prime}+1$ of $\mathcal{A}$. Finally, by construction, gate $v^{(2\ell^{\prime}+1)}_{i,n+j}$ for $j\in[n,(\ell^{\prime}+1)n]$ gets and outputs the wire from gate $v^{(2\ell^{\prime})}_{i,j}$ and thus corresponds to the input of node $i$ and all messages it received in rounds 1 to $\ell^{\prime}$, as needed. ∎